OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. 5 + yi = x + 3i

OpenStudy (unklerhaukus):

\[5 + yi = x + 3i\] \[5=x;\qquad y=3\]

OpenStudy (anonymous):

What ^ did is basically bring the x and y terms to one side. \[5 + yi = x + 3i\]\[x - yi = 5 - 3i\]Then separate the real and imaginary portions since they cannot mix. \[x = 5\]\[-yi = -3i\]\[y = 3\]

OpenStudy (anonymous):

oh okay thank you for explaining it to me ^_^"

OpenStudy (anonymous):

np :)